OpenStudy (anonymous):

Solve equation. If there is no solution or an identity write 0 or identity. 15-3(y+1)=14-5y

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

15-3y-3=14-5y 12-3y =14-5y -3y+5y =14-12 2y=2 y=2/2 y =1

OpenStudy (anonymous):

TY